# Enhancing Event Extraction from Short Stories through Contextualized Prompts

### v1

@@ 3.3 Evaluation Metrices
Table 1: Top 15 words that are annotated as event triggers. Its count and event rate is also shown. Event rate is the percentage of labeled events with respect to total occurrence of that word in the corpus
4 Event Detection and Classification
In this section, the methods employed for event detection and classification are discussed. Two different approaches for the task are defined. These methods are outlined and elaborated in the subsequent subsections.
@@ 4.1 Neural Methods
Previous research has shown that neural models are good at detecting events. These models use the distributional information encoded in word embeddings. In this work, different variants of these models are analyzed, all of which take a sequence labeling approach, providing labels to tokens to indicate their event state. Considering $\mathcal{D}$ as the entire dataset & $E$ as a set of event types, we formulate the task over a document $\mathcal{D}_{i}\in\mathcal{D}$, where an instance of the document $\mathcal{D}_{i}$ is defined as $(X,t,e_{i})\in\mathcal{D}_{i}$, where $X$ represents the sentence, $t$ represents the trigger word and $e_{i}$ represents the event type associated with it. The task targeted in this section is to extract the event trigger $t$ and its type $e_{i}$ from $X$ given that $t\subseteq X$ $e_{i}\in E$. To achieve this, several neural models were explored in the literature. We chose to employ these models in our current work. The task of event trigger detection is framed as a sequence labeling task. The overall framework is shown in Figure 3. The neural model in this framework consists of the following three modules – (a) Embedding Generation, (b) Sequence to Vector and (c) Dense Layer.
Figure 3: Block diagram illustrating the event detection framework.
As shown in Figure 3, the ED task involves 3 inner modules - Embedding Generation, Sequence to Vector, and a Dense Layer which will learn the function to predict one among the 7 output classes.
@@
In the above statement, London plays an important role in classifying came as a separate type of event class, when compared with fighting. In the latter came describes a CONFLICT event class whereas in the first case, it was a MOVEMENT event class. Hence, it is clear that for better prediction we need forward as well as backward information. Thereby we can conclude that BiLSTM will serve as a great basic model for the detection and classification of an event. Further modification of the BiLSTM is done as described below so as to capture more information. However, most of them have not improved the performance of the basic model significantly as will be clear in the results section.
BiLSTM with Sentence CNN Several works have used a CNN on the sentence level nguyen2015event for the task of event detection. When we are predicting the event status of a token $w_{i}$ in a sentence that consists of $n$ words such that $w_{i}\in W$ and $W=[w_{1},w_{2},w_{3},...,w_{i},...,w_{n}]$, each CNN (1-dimensional CNN) convolves over the entire sequence $W$ along with the positional embeddings $P=\{p_{1},….,p_{n}\}$ which encode the distance between each token $w_{j}$ and the target token $w_{i}$. We have adopted the work of Nguyen et. al. nguyen2015event where the output of the CNN is then passed to a max-pooling phase to yield a representation $c_{i}$ for the token $w_{i}$ that is concatenated to the BiLSTM output $o_{i}$ for the same time step before passing it to the dense layer for prediction.
There are two CNN layers of 100 filters one considering bi-gram token representation (kernel=2), the other representing tri-gram representation(kernel=3). The positional arguments are encoded between the target token at position $i$ and the token at position $j$ using signed bucketing (±1,2,3,4,5,($\pm$ 1,2,3,4,5,( ± 1 , 2 , 3 , 4 , 5 ,6-10,11-$,>20)$. Each bucket corresponds to a discrete choice of position with its 5-dimensional embedding (as was done in the past work). Time-distributed layers signify that weights are not being shared across the timestamps so that each timestamp/word can have its separate prediction.
BiLSTM with Subword CNN Subword character CNN captures meaningful representations of out-of-vocabulary words for learned embeddings. The architecture for character CNN considers each word to be represented as the output of a CNN with 100 filters, i.e., 25 filters scoped over character bi-grams (kernel=2), tri-grams (kernel=3), 4-grams (kernel=4) and 5-grams (kernel=5). The output from the convolution layer is then fed to the max pooling layer, hence getting a 100-dimensional character representation of a word. This representation is then concatenated with the skip-gram word embedding for the token, finally giving it as input to the sequential model for the respective timestamp.
BiLSTM with Document Context The main problem with BiLSTM is that the only information available is from the set, which may not always be enough to say exactly whether a given token represents an event. Therefore, to improve the performance of BiLSTM, we integrated the document context in addition to the sentence context already captured by the model. Based on previous research involving the global context liao2010using, we can conclude that accurate prediction of complex Realis events requires a lot of document context across pages or documents. Therefore, in this experiment, we defined sequences as entire documents rather than as single sentences. Training a model like this involves a lot of parameters and takes more time to train the model, but it did not provide a significant improvement in the results obtained by the previous vanilla model.
BiLSTM with CRF In this architecture, the BiLSTM is used to extract features from the input sequence, which are then used as inputs to the CRF [wang2018bidirectional]. Using a basic BiLSTM followed by a classifier means that each classification decision is made independently of the others. In contrast, a linear-chain CRF represents dependencies between labels by using a table that contains transition scores for all label pairs, explicitly modeling these dependencies. The CRF layer considers both the output scores from the BiLSTM and the transition matrix to calculate the probability of the entire sequence of labels. During training, the model learns the parameters of both the BiLSTM (for feature extraction) and the CRF (for label dependencies) jointly, typically using the negative log-likelihood as the loss function. During inference, the Viterbi algorithm is often used to find the most likely sequence of labels based on the output scores from the BiLSTM and the learned transition matrix from the CRF. This combination of BiLSTM and CRF is effective for sequence labeling tasks like event detection because it allows the model to capture complex dependencies between words in the input sequence and output labels, leading to more accurate predictions.
@@
BERT fine-tuning Fine-tuning BERT involves customizing a pre-trained BERT model for a specific task or domain by adjusting its parameters using a small amount of labelled data. For instance, in our scenario, fine-tuning involves using a dataset that includes texts paired with their corresponding event class labels for event classification with BERT. This usually involves adding a task-specific layer on top of the BERT encoder and training the entire model together, using a suitable loss function and optimizer.
@@
The training for the models was conducted on 120 short stories, with a validation set and testing set consisting of consisting of 20 and 60 short stories respectively. The optimizer that we have used is ADAM [DBLP:journals/corr/KingmaB14] with Binary Cross-Entropy as the loss function for event trigger detection, and Categorical Cross-Entropy for predicting the categories of the event triggers detected. We have used ReLU [agarap2018deep] as the activation function for 1-D convolutional layers. The training of each model was done for 1000 epochs. The batch size was 16, and it took about 2 hours to converge.
@@
The above models primarily focused on trigger detection. We adapted these models for multi-class classification. In the field of event extraction, prompt-based generative approaches have recently emerged that offer greater flexibility than earlier classification-based methods.
4.2 Prompt Based
Prompt-based learning represents a novel approach within the domain of pre-trained language models [liu2023pre]. Diverging from the traditional pre-training and fine-tuning paradigm, prompt-based techniques transform downstream tasks into formats that align more closely with the model’s pre-training objectives. Contextualized prompts can be used in conjunction with pre-trained language models to fine-tune the model for event extraction tasks. The prompts help the model understand the specific requirements of the task and improve its performance by providing the model with relevant context for understanding and extracting events from text. These prompts typically consist of a combination of text and placeholders that guide the model on what information to extract and how to structure the output.
4.2.1 Problem Formulation
4.2.2 Proposed Architecture
Existing works try to model the event detection task by first encoding the input words, taking out its embedding, and then learning a function that can map the input vector to the corresponding event type. This increases the number of parameters required and also the amount of data that we have to provide to learn the mapping function. These formulations are highly resource intensive and prone to biasing one category with more instances in the dataset.
Prompt-based approaches try to convert these types of problems into classification and generation tasks. However, this creates a problem with efficiency since the model must list every candidate we wish to extract from the text. cui-etal-2021-template proposed an architecture for named entity recognition, which involved prompts like - [candidate span] is [entity type/not a] entity. This was inefficient since the models had to initially fill the first slot with potential entities before checking the second slot’s output from the language model. In the proposed architecture, we compressed the event detection and classification in a single iteration.
  1. (1)
    Prompt Creation: Considering the context $X_{j}$ is provided as the input, if the queries that we want are $Q=[q_{1},q_{2},...,q_{n}]$ a joint prompt was created by bringing all the queries together. In the case of event extraction, only one query is sufficient as for all the sentences, we want to find the triggers $t$ corresponding to the same set of event $E$. In the case of event classification, all the event classes are provided as input through prompts, which is depicted in Table 2. This provides context to the model, indicating which category the word or phrase under consideration falls into, based on the provided event types as a contextualized prompt.
    $P=f_{prompt}(Q)$ (1)
  2. (2)
    Prompt Decoding: Considering a pre-trained language model $\mathcal{L}$, prompt $P$, and the context $X$, we can identify a query-distinct (answering) span selector as per Equation 2.
    $\theta_{q_{n}}=O_{\mathcal{L}}(q_{n};P,X)$ (2)
    where $q$ is the query for which the prompt is formulated and $O_{\mathcal{L}}$ is the output obtained from the pre-trained language model.
  3. (3)
    Span selection: To find the span for the starting token and ending token for the query $q$, we use two selectors for the beginning and completion of tokens from the context as given in the Equation 3.
    $<s,e>_{q_{n}}=f_{span}(X_{j};\theta_{q_{n}})$ (3)
This way of formulating event extraction is better because it can be extended for other types of information extraction tasks.
4.2.3 Prompt Creation
We create a single prompt for the types of events that we want to extract from the dataset $\mathcal{D}$. The presence of event types in the prompt is referred to as the slots for the events. We don’t allow multiple events of the same type to occur in a context, because the dataset $\mathcal{D}$ doesn’t have multiple events of the same type in the same context. In our work, we have used concatenation to create the prompt where we concatenate all the event types that we want to extract from $\mathcal{D}$ as shown in the Table 2. This might not capture the interaction among different events present in the context, however, in the dataset under consideration it is assumed that more than one event of the same type is not present in the same context.
Table 2: Prompts that are used by the prompt creator function
4.2.4 Learning prompt-based span selection
Considering the context $X$ and the prompt $P$ as the input, we first aim to create a learnable span selector for each event $e_{k}\in E$. We choose BART [lewis-etal-2020-bart], a pre-trained language model built on a standard transformer that combines both an encoder and a decoder which we represent here as $\mathcal{L}=[\mathcal{L}_{enc},\mathcal{L}_{dec}]$. We do not concatenate the prompt $P$ and the context $X$, instead, we give the context $X$ to the BART-Encoder $\mathcal{L}_{enc}$ and prompt $P$ to BART-Decoder $\mathcal{L}_{dec}$. $X$ and $P$ would interact mutually in the cross-attention layers of $\mathcal{L}_{dec}$.
$H^{(enc)}_{X}=\mathcal{L}_{enc}(X)$ (4)
$H_{X}=\mathcal{L}_{dec}(H^{(enc)}_{X};H^{(enc)}_{X})$ (5)
$H_{P}=\mathcal{L}_{dec}(P;H^{(enc)}_{X})$ (6)
where $H_{X}$ denotes the context representation and $H_{P}$ denotes to context-oriented prompt representation. If we want to extract the context-oriented prompt representation for the $k^{th}$ event we have to mean-pool its corresponding representation in $H_{P}$ and obtain the event feature $\lambda_{k}\in R^{h}$ where $h$ denotes the dimension of the last hidden layer of $\mathcal{L}_{dec}$.
$H_{P_{k}}=({H_{P}}_{ij})~{}\forall_{j}\in[P^{st}_{k},P^{ed}_{k}]$ (7)
[…]
where $P^{st}_{k}$ is start and $P^{ed}_{k}$ is the end of $k^{th}$ slot in prompt $P$. After extracting the representation for the $k^{th}$ event, we can utilize the same for learning the starting and ending position for that event in the context representation $H_{X}$. To do the same we use a span selector $\theta=[w^{(start)},w^{(end)}]\in R^{h\times 2}$ with random initialization, as the learnable parameters. So, for an event feature $\lambda_{k}$ we have:
$\lambda^{(start)}_{k}=\lambda_{k}\circ w^{(start)}\in R^{h}$ (9)
$\lambda^{(end)}_{k}=\lambda_{k}\circ w^{(end)}\in R^{h}$ (10)
$logit^{(start)}_{k}=\lambda^{(start)}_{k}\times H_{X}\in R^{L}$ (11)
$logit^{(end)}_{k}=\lambda^{(end)}_{k}\times H_{X}\in R^{L}$ (12)
where $logit^{(start)}_{k}$ and $logit^{(end)}_{k}$ represent the start and the end token distribution over the context tokens for $e_{k}$, and $L$ indicates the context length, we calculate the probabilities where the start or the end positions are located by converting the distribution into a stochastic distribution using Softmax.
$p^{(start)}_{k}=SoftMax(logit^{(start)}_{k})\in R^{L}$ (13)
$p^{(end)}_{k}=SoftMax(logit^{(end)}_{k})\in R^{L}$ (14)
$L_{k}(X)=-(\log p^{(start)}_{k}(st_{k})+\log p^{(end)}_{k}(ed_{k}))$ (15)
$L=\sum_{X\in D}\sum_{k}L_{k}(X)$ (16)
where $D\in[D_{1},D_{2},...D_{n}]$ which makes up the entire dataset $\mathcal{D}$ and $k\in[1,l]$ denotes the event $e_{k}$.
@@ 4.2.4 Learning prompt-based span selection
4.2.5 Inference
First, we define a set of spans for the events as $\mathcal{S}=\{(i,j)|(i,j)\in L^{2},0<j-i\leq\alpha\}\cup\{(0,0)\}$. It consists of all the spans that are less than the threshold $\alpha$ and a special span $\{(0,0)\}$ to mark it as an invalid event. Our model extracts the event trigger related to the span selector $\theta_{k}$ by iterating and scoring all candidates as shown
$score_{k}(i,j)=logit^{(start)}_{k}(i)+logit^{(end)}_{k}(j)$ (17)
[…]
We train event extraction and event argument Extraction separately in the training phase. After the training of Event Extraction is complete, we tag the sentences with the extracted events and then train the Event Argument Extraction using the predictions from the Event Extraction phase. For the testing phase, we tested the Event Extraction phase separately to compare it with the previous work that we had done. We show the class-wise Precision, Recall, and F1 score for better comparison. In Table 3, we list the hyperparameters that have been used for training the model.
@@ 4.2.5 Inference
5 Results and Discussion
5.1 Dataset Statistics
The proposed annotation guidelines adapts the problems and rules addressed by ACE [ace] for the chosen data. Only the realis events, i.e., the events that are actually happening, with their event classes are considered  [doddington2004automatic]. For capturing realis events, four aspects - Polarity, Tense, Genericity, and Modality have been selected [aguilar-etal-2014-comparison][song2015light][lit].
Sl.No Statistics Count
1. Total tokens in the dataset 174,559
2. Total unique tokens in the dataset 10,484
3. Total sentences in the dataset 10,259
4. Average tokens per story 873
5. Average sentences per story 51
6. Average tokens per sentence 17
7. Total events in the dataset 11,272
8. Average events per story 56
Table 4: Statistics of the annotated dataset (200 short stories)
We annotate 200 stories based on the proposed guideline. An equal number of 25 stories have been considered from each of the sources to avoid biases related to a particular narrative. The web-based BRAT annotation tool [a18] is used for annotation. The number of words in the stories ranges between 500 to 1500. Two annotators make all the annotations based on the proposed guideline. The distribution of the events from different short stories is shown in Figure 7. The figure shows that the events depicting the cognitive mental state are the least in number. This study presents an annotated dataset consisting of realis events from Indian short stories with 11,272 overall events. Table 4 shows a detailed breakdown of the dataset statistics.
Event Types Total Count Train Dev Test
COMMUNICATION 3456 2052 447 957
GENERAL-ACTIVITY 2951 1837 325 789
MOVEMENT 2192 1329 294 569
COGNITIVE-MENTAL-STATE 1528 849 178 501
LIFE-EVENT 535 314 60 161
OTHERS 490 295 55 140
CONFLICT 120 62 30 28
Total 11,272 6,738 1,389 3,145
Table 5: Number of different events present in the annotated dataset
Figure 6: Distribution of different events in the dataset
Table 5 presents the total count of events within the dataset, along with the distribution of events across the train, dev, and test splits. Figure 6 represents the distribution of the events in the dataset. Table 1 represents the top 15 words tagged as an event in the annotated dataset. Figure 2 presents the word cloud of the entire dataset and for trigger words only. It is clear that trigger words with the highest frequencies are in the majority of the entire dataset. This can also be correlated with Table 1. Words like “said,” “asked,” “went,” “came,” “told,” etc., are evident in the dataset. Two annotators, who are co-authors of this work, performed the annotation task. 20 random stories were chosen, and both the annotators annotated the stories in accordance with the provided annotation guidelines. An inter-annotator agreement (IAA) of 83.1% was achieved. The IAA serves as a meaningful indicator of the clarity of the annotation guidelines, the consistency in the annotators’ comprehension of these guidelines, and the overall feasibility of the annotation task. The remaining 180 stories were annotated by a single annotator, resulting in a total of 200 annotated stories.
The event trigger detection results obtained from the baseline models and prompt-based model are shown in Table 6. It is observed that the proposed model EVENT PROMPT yields an F1 score of 89.9% which is higher than the baseline models. In a previous study, from which we adopted baseline models [sims2019literary], the F1 scores achieved for various models utilized in event detection from literary texts were lower than those obtained for short stories. In the case of event classification, it is observed that except for the class MOVEMENT, the other classes classified by EVENT PROMPT with an F1 score are better than the other baseline models. Out of all the classes, the COMMUNICATION class yields the highest F1 score, and the class OTHERS have the lowest. This might be due to the greater number of events for the class COMMUNUCATION as compared to the other classes Table 5. We noted a significant imbalance in the distribution of event types within the dataset. This factor can be attributed to the fact that in children’s short stories, the teachings, morals, discipline, etc., taught to the children are narrated through verbal communication between the subjects of the stories. The target audience for the short stories is children, and the aim of these stories is to nurture young minds with positive and constructive ideas while steering them away from harmful or negative thoughts. This fact is clearly demonstrated by the data presented in Table 5. It shows a notable disparity in the distribution of event types, with the number of events classified as CONFLICT being significantly lower compared to the maximum number of events falling under the COMMUNICATION category. Table 1 clearly illustrates that most events depicted in children’s literature are based on reality. These narratives are grounded in real-life experiences, facts, and occurrences, emphasizing practical teachings and tangible lessons. Conversely, instances of purely imaginative or fantastical ideas are comparatively scarce within this genre. This indicates a deliberate emphasis on imparting realistic knowledge and values to young readers, enriching their understanding of the world around them. Figure 5 illustrate an example demonstrating the detailed response provided by our proposed model. The input prompt, as well as the output from the model, are shown.
METHOD PRECISION RECALL F1
LSTM 87.7 85.3 86.8
BiLSTM 86.0 88.8 87.7
+Document Context 87.6 66.5 75.9
+Sentence CNN 86.0 85.9 86.3
+Char CNN 86.5 88.5 87.8
+CRF 87.0 92.2 89.5
BERT (fine-tuned) 86.2 93.2 89.7
EVENT PROMPT 87.5 92.4 89.9
Table 6: Event trigger detection
@@ 8. Average events per story 56
In Table 7, we can observe that the best-performing model from the previously discussed methods is outperformed by the prompt-based approach, especially for the categories with very few occurrences (e.g. CONFLICT class). This is mainly because of the presence of better context features before the extraction of the span. Better analysis can be done by observing the confusion matrix shown in Figure 8 for the EVENT PROMPT model. The maximum misclassification can be seen in GENERAL-ACTIVITY and OTHERS class. Mostly, the class GENERAL-ACTIVITY was misclassified from the other classes as the scope of this class is greater compared to the others. The annotators get confused when marking this event class.

### v2

@@ 8. Average events per story 56
Table 3: Top 15 words that are annotated as event triggers. Its count and event rate are also shown. Event rate is the percentage of labelled events with respect to the total occurrence of that word in the corpus
3.4 Evaluation Metrices
We used the identical evaluation criteria as previous studies [Wadden et al., 2019], which presented the Precision (P), Recall (R), and F1 scores for event detection. We defined TP as a true positive, FN as a false negative, and FP as a false positive.
$P=\frac{TP}{TP+FP}\quad$
$R=\frac{TP}{TP+FN}\quad$
$F1=2*\frac{P*R}{P+R}\quad$
4 Contextualized Prompts
Prompt-based learning represents a novel approach within the domain of pre-trained language models [Liu et al., 2023]. Diverging from the traditional pre-training and fine-tuning paradigm, prompt-based techniques transform downstream tasks into formats that align more closely with the model’s pre-training objectives. Contextualized prompts can be used in conjunction with pre-trained language models to fine-tune the model for event extraction tasks.
4.1 Problem Formulation
4.2 Proposed Architecture
4.2.1 Prompt Creation
We create a single prompt for the types of events that we want to extract from the dataset $\mathcal{D}$. The presence of event types in the prompt is referred to as the slots for the events. We don’t allow multiple events of the same type to occur in a context, because the dataset $\mathcal{D}$ doesn’t have multiple events of the same type in the same context. In our work, we have used concatenation to create the prompt where we concatenate all the event types that we want to extract from $\mathcal{D}$ as shown in the Table 4.
Table 4: Prompts that are used by the prompt creator function
4.2.2 Learning prompt-based span selection
Considering the context $X\in\mathbb{R}^{t_{x}\times d_{e}}$ where $t_{x}$ is the length of input tokens and $d_{e}$ is the output dimension of feature embedding, along with prompt $P$ of length $t_{p}$ as the conditioning input, we first aim to create a learnable span selector for each event $e_{k}\in E$. We choose BART [Lewis et al., 2020], a pre-trained language model built on a standard transformer that combines both an encoder and a decoder which we represent here as $\mathcal{L}=[\mathcal{L}_{enc},\mathcal{L}_{dec}]$. We do not concatenate the prompt $P$ and the context $X$, instead, we give the context $X$ to the BART-Encoder $\mathcal{L}_{enc}$ and prompt $P$ to BART-Decoder $\mathcal{L}_{dec}$. $X$ and $P$ would interact mutually in the cross-attention layers of $\mathcal{L}_{dec}$.
$H^{(enc)}_{X}=\mathcal{L}_{enc}(X)\in\mathbb{R}^{t_{x}\times d_{enc}}$ (1)
$H_{X}=\mathcal{L}_{dec}(H^{(enc)}_{X};H^{(enc)}_{X})\in\mathbb{R}^{t_{x}\times h}$ (2)
$H_{P}=\mathcal{L}_{dec}(P;H^{(enc)}_{X})\in\mathbb{R}^{t_{p}\times h}$ (3)
where $H_{X}$ denotes the context representation and $H_{P}$ denotes to context-oriented prompt representation. If we want to extract the context-oriented prompt representation for the $k^{th}$ event we have to mean-pool its corresponding representation in $H_{P}$ and obtain the event feature $\lambda_{k}\in\mathbb{R}^{h}$ where $h$ denotes the dimension of the last hidden layer of $\mathcal{L}_{dec}$.
$H_{P_{k}}=({H_{P}}_{ij})~\forall_{j}\in[P^{st}_{k},P^{ed}_{k}]\mid H_{P_{k}}\in\mathbb{R}^{end(t_{p_{k}})-start(t_{p_{k}})\times h}$ (4)
$\lambda_{k}=\operatorname*{Mean\,Pool}(H_{P_{k}})\in\mathbb{R}^{h}$ (5)
where $P^{st}_{k}$ is start and $P^{ed}_{k}$ is the end of $k^{th}$ slot in prompt $P$. After extracting the representation for the $k^{th}$ event, we can utilize the same for learning the starting and ending position for that event in the context representation $H_{X}$. To do the same we use a span selector $\theta=[w^{(start)},w^{(end)}]\in\mathbb{R}^{2h}$ with random initialization, as the learnable parameters. So, for an event feature $\lambda_{k}$ we have:
$\lambda^{(start)}_{k}=\lambda_{k}\circ w^{(start)}\in\mathbb{R}^{h}$ (7)
$\lambda^{(end)}_{k}=\lambda_{k}\circ w^{(end)}\in\mathbb{R}^{h}$ (8)
$logit^{(start)}_{k}=\lambda^{(start)}_{k}\times({H_{X}})^{T}\in\mathbb{R}^{t_{x}}$ (9)
$logit^{(end)}_{k}=\lambda^{(end)}_{k}\times({H_{X}})^{T}\in\mathbb{R}^{t_{x}}$ (10)
where $logit^{(start)}_{k}$ and $logit^{(end)}_{k}$ represent the start and the end token distribution over the context tokens for $e_{k}$, and $t_{x}$ indicates the context token length, we calculate the probabilities where the start or the end positions are located by converting the distribution into a stochastic distribution using Softmax.
$p^{(start)}_{k}=SoftMax(logit^{(start)}_{k})\in\mathbb{R}^{t_{x}}$ (11)
$p^{(end)}_{k}=SoftMax(logit^{(end)}_{k})\in\mathbb{R}^{t_{x}}$ (12)
$L_{k}(X)=-(\log p^{(start)}_{k}(st_{k})+\log p^{(end)}_{k}(ed_{k}))$ (13)
$L=\sum_{X\in D}\sum_{k}L_{k}(X)$ (14)
where $D\in[D_{1},D_{2},...D_{n}]$ which makes up the entire dataset $\mathcal{D}$ and $k\in[1,l]$ denotes the event $e_{k}$ and $l$ is the number of event categories in prompt.
4.2.3 Inference
First, we define a set of spans for the events as $\mathcal{S}=\{(i,j)|(i,j)\in{t_{x}}^{2},0<j-i\leq\alpha\}\cup\{(0,0)\}$. It consists of all the spans that are less than the threshold $\alpha$ and a special span $\{(0,0)\}$ to mark it as an invalid event. Our model extracts the event trigger related to the span selector $\theta_{k}$ by iterating and scoring all candidates as shown
$score_{(k)}(i,j)=logit^{(start)}_{k}(i)+logit^{(end)}_{k}(j)$ (15)
$(\hat{st}_{k},\hat{ed}_{k})=\operatorname*{arg\,max}_{(i,j)\in\mathcal{S}}score_{(k)}(i,j)$ (16)
5 Experiments
Previous research has shown that neural models are good at detecting events Sims et al. [2019b]. These models use the distributional information encoded in word embeddings. In this work, different variants of these models are analyzed, all of which take a sequence labeling approach, providing labels to tokens to indicate their event state. To achieve this, several neural models were explored in the literature. We chose to employ these models in our current work. The task of event trigger detection is framed as a sequence labelling task. The overall framework is shown in Figure 6. The neural model in this framework consists of the following three modules – (a) Embedding Generation, (b) Sequence to Vector and (c) Dense Layer, which will learn the function to predict one among the 7 output classes.
Figure 6: Block diagram illustrating the event detection framework.
@@ 5.1 Neural Methods
In the above statement, London plays an important role in classifying came as a separate type of event class, when compared with fighting. In the latter came describes a CONFLICT event class whereas in the first case, it was a MOVEMENT event class. Hence, it is clear that for better prediction we need forward as well as backward information.
BiLSTM with Sentence CNN Several works have used a CNN on the sentence level for the task of event detection [Nguyen and Grishman, 2015]. We have employed the model used by Sims et al. [2019b] for event detection and modified the same for event classification.
BiLSTM with Subword CNN Subword character CNN captures meaningful representations of out-of-vocabulary words for learned embeddings.
BiLSTM with Document Context The main problem with BiLSTM is that the only information available is from the set, which may not always be enough to say exactly whether a given token represents an event. Therefore, to improve the performance of BiLSTM, we integrated the document context in addition to the sentence context already captured by the model. We have utilized the same model that Sims et al. [2019b] have used for this purpose. Necessary modifications were done for the task of event classification.
BiLSTM with CRF In this architecture, the BiLSTM is used to extract features from the input sequence, which are then used as inputs to the CRF [Wang et al., 2018].
@@
BERT fine-tuning Fine-tuning BERT involves customizing a pre-trained BERT model for a specific task or domain by adjusting its parameters using a small amount of labelled data. For instance, in our scenario, fine-tuning involves using a dataset that includes texts paired with their corresponding event class labels for event classification with BERT. This involves adding a task-specific layer on top of the BERT encoder and training the entire model together, using a suitable loss function and optimizer.
EVENT PROMPT
@@
The training for the models was conducted on 120 short stories, with a validation set and testing set consisting of consisting of 20 and 60 short stories respectively. The optimizer that we have used is ADAM [Kingma and Ba, 2015] with Binary Cross-Entropy as the loss function for event trigger detection, and Categorical Cross-Entropy for predicting the categories of the event triggers detected. We have used ReLU [Agarap, 2018] as the activation function. The training of each model was done for 100 epochs. The batch size was 16, and it took about 2 hours to converge.
@@
The above models primarily focused on trigger detection. We adapted these models for multi-class classification for the task of event classification. In the field of event extraction, prompt-based generative approaches have recently emerged that offer greater flexibility than earlier classification-based methods.
We show the class-wise Precision, Recall, and F1 score for better comparison. In Table 5, we list the hyperparameters that have been used for training the model.
@@
6 Results and Discussion
The event trigger detection results obtained from the baseline models and prompt-based model are shown in Table 6. It is observed that the proposed model EVENT PROMPT yields an F1 score of 89.9% which is higher than the baseline models. In a previous study, from which we adopted baseline models [Sims et al., 2019b], the F1 scores achieved for various models utilized in event detection from literary texts were lower than those obtained for short stories. In the case of event classification, it is observed that except for the class MOVEMENT, the other classes classified by EVENT PROMPT with an F1 score are better than the other baseline models. Out of all the classes, the COMMUNICATION class yields the highest F1 score, and the class OTHERS have the lowest. This might be due to the greater number of events for the class COMMUNUCATION as compared to the other classes as shown in Table 2. We noted a significant imbalance in the distribution of event types within the dataset.
METHOD PRECISION RECALL F1
EVENT PROMPT 85.35 87.23 86.28
Table 6: Result of Event trigger detection
@@ 6 Results and Discussion
This factor can be attributed to the fact that in children’s short stories, the teachings, morals, discipline, etc., taught to the children are narrated through verbal communication between the subjects of the stories. The target audience for the short stories is children, and the aim of these stories is to nurture young minds with positive and constructive ideas while steering them away from harmful or negative thoughts. This fact is clearly demonstrated by the data presented in Table 2. It shows a notable disparity in the distribution of event types, with the number of events classified as CONFLICT being significantly lower compared to the maximum number of events falling under the COMMUNICATION category. Table 3 clearly illustrates that most events depicted in children’s literature are based on reality. These narratives are grounded in real-life experiences, facts, and occurrences, emphasizing practical teachings and tangible lessons. Conversely, instances of purely imaginative or fantastical ideas are comparatively scarce within this genre. This indicates a deliberate emphasis on imparting realistic knowledge and values to young readers, enriching their understanding of the world around them. Figure 7 illustrate an example demonstrating the detailed response provided by our proposed model. The input prompt, as well as the output from the model, are shown.
In Table 7, we can observe that the best-performing model from the previously discussed methods is outperformed by the prompt-based approach, especially for the categories with very few occurrences (e.g. CONFLICT class). This is mainly because of the presence of better context features before the extraction of the span. Better analysis can be done by observing the confusion matrix shown in Figure 8 for the EVENT PROMPT model. The maximum misclassification can be seen in GENERAL-ACTIVITY and OTHERS class. Mostly, the class OTHERS was misclassified from the other classes as the scope of this class is greater compared to the other classes. The annotators get confused when marking this event class.